# Enhancing Speech Large Language Models with Prompt-Aware Mixture of Audio Encoders

### v1

@@ Abstract
Connecting audio encoders with large language models (LLMs) allows the LLM to perform various audio understanding tasks, such as automatic speech recognition (ASR) and audio captioning (AC). Most research focuses on training an adapter layer to generate a unified audio feature for the LLM. However, different tasks may require distinct features that emphasize either semantic or acoustic aspects, making task-specific audio features more desirable. In this paper, we propose Prompt-aware Mixture (PaM) to enhance the Speech LLM that uses multiple audio encoders. Our approach involves using different experts to extract different features based on the prompt that indicates different tasks. Experiments demonstrate that with PaM, only one Speech LLM surpasses the best performances achieved by all single-encoder Speech LLMs on ASR, Speaker Number Verification, and AC tasks. PaM also outperforms other feature fusion baselines, such as concatenation and averaging.
@@
Weiqiao Shan1, Yuang Li2thanks: Equal contribution., Yuhao Zhang3, Yingfeng Luo1, Chen Xu4, Xiaofeng Zhao2, Long Meng1, Yunfei Lu2, Min Zhang2, Hao Yang2, Tong Xiao1,5thanks: Corresponding author, Jingbo Zhu1,5, 1School of Computer Science and Engineering, Northeastern University, Shenyang, China, 2Huawei Translation Services Center, Beijing, China, 3The Chinese University of Hong Kong, Shenzhen, China, 4College of Computer Science and Technology, Harbin Engineering University, Harbin, China, 5NiuTrans Research, Shenyang, China
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance across various natural language processing tasks, including question answering, machine translation, and summarization OpenAI (2023). In recent work, there has been a growing focus on merging speech encoders with LLMs, so that the LLM can understand the spoken content without the need for explicit transcription, promoting tasks such as direct speech translation Chen et al. (2024) and named entity recognition from speech Li et al. (2024). Much of this work leverages adapter layers to downsample and map speech features into the LLM’s embedding space. These adapter layers include various forms, such as attention layers Yu et al. (2024), adaptive CTC downsamplers Ling et al. (2023), and convolutional layers Fathullah et al. (2023). Beyond semantic understanding tasks, Speech LLMs have been extended to encompass a broader range of applications, including audio event detection and audio captioning Chu et al. (2024).
Figure 1: ASR and Audio Caption tasks favor different encoders and layers of features. The dotted lines indicate the average (AVG) importance of different encoders. The bar chart represents fine-grained layer importance.
@@
Consequently, researchers have proposed using multiple encoders to extract more robust features. For instance, WavLLM Hu et al. (2024) employs both the WavLM Chen et al. (2022) and the Whisper Radford et al. (2022) encoder, while SALMONN Tang et al. (2024) integrates the Whisper encoder and the BEATs Chen et al. (2023). However, these approaches consider all encoders equally important, and merge the features from different encoders based on simple concatenation method across all tasks. Moreover, MoWE Zhang et al. (2024) employs a strong encoder and multiple weaker encoders via the Mixture of experts (MoE) approach. However, MoWE only utilizes the input audio to control the gating mechanism, without incorporating task-specific information in prompts.
In this paper, we introduce Prompt-aware Mixture (PaM), a novel MoE method for merging multiple encoders to enhance Speech LLMs, our approach integrates a prompt-aware gating mechanism, emphasizes feature fusion, and considers the relative importance of each encoder for different tasks, aiming to improve all downstream performance. PaM employs three distinct audio encoders: the Whisper encoder Radford et al. (2022), WavLM Chen et al. (2022), and Wav2Vec2 Baevski et al. (2020). We train a set of experts for prompt-aware feature fusion, comprising one shared expert and four task-specific experts. On each task, an expert learns the optimal weights for each encoder and its respective layers, and subsequently maps the resulting features to the embedding space of the Qwen2.5 model Team (2024). The embedding of the prompt is utilized to determine the appropriate routing. Notably, in PaM, the routing guides the selection of the fusion parameters rather than the choice of the encoder. Experiments are conducted across three tasks: ASR, speaker number verification, and AC. On all datasets, including LibriSpeech Panayotov et al. (2015), AMI Kraaij et al. (2005), AIR-Bench (speaker number verification) Yang et al. (2024), and AudioCaps Kim et al. (2019), PaM consistently enhances performance in comparison to the best-performing single-encoder Speech LLM. The relative improvements are 15%, 25%, 3.4%, and 7.6%, respectively, based on the corresponding evaluation metrics. Additionally, we observed that without PaM, simple concatenation and averaging are generally effective for AC tasks; however, they result in performance declines for ASR. Our contributions can be summarized as follows:
@@
  • We introduce PaM, a specialized MoE method that incorporates a prompt-aware gating mechanism to assign distinct weights to each encoder and their layers based on the task.
@@
Figure 2: The architecture of the proposed PaM method. The output feature of the prompt guides the routing of the MoE adapter where we use a fixed shared expert and a single routed expert. For each expert, the last hidden states of all encoders are concatenated with a set of fused hidden states derived from a fusion weight matrix. Subsequently, an FFN is applied to align with the dimensions of the LLM.
@@ 2.1 Overall Architecture
The architecture of the proposed PaM method is depicted in Figure 2. As described in Equation 1, the LLM accepts the text prompt $\mathbf{X}_{\text{prompt}}$, which includes task-related information, along with the speech features $\mathbf{H}_{\text{audio}}$ as input, and subsequently generates the response $\mathbf{Y}$.
@@
To obtain $\mathbf{H}_{\text{audio}}$, we employ three encoders: the Whisper encoder, WavLM, and Wav2Vec2. For each encoder, the hidden states are initially processed by a feed-forward network (FFN) as described in Equation 2, resulting in the feature of each encoder, denoted as $\mathbf{H}_{i}$.
@@
$+\sum_{j=1}^{N}G_{j}(\mathbf{X}_{\text{prompt}})\times\text{Expert}_{j}(\mathbf{H}_{\{1,2,3\}})$ (3)
@@ 2.2 Multi-layer Fusion
Different encoders exhibit distinct strengths. For example, WavLM is excellent at extracting speaker information, while Wav2Vec2 excels in capturing semantic content. The Whisper encoder, trained on a vast amount of data, provides superior features for AC and ASR in noisy environments. Additionally, features from different layers contain varying levels of information. Deeper layers hold high-level semantic information, whereas lower layers may contain fine-grained acoustic details. Thus, for feature fusion, we consider features from all layers of all three encoders. Specifically, for the feature $\mathbf{H}_{i}$ from a single encoder, it includes hidden states from all $L_{i}$ Transformer Vaswani et al. (2017) layers as well as $\mathbf{h}_{i}^{0}$, the hidden states following the convolutional layers (Equation 4).
$\mathbf{H}_{i}=\{\mathbf{h}_{i}^{0},\mathbf{h}_{i}^{1},...\mathbf{h}_{i}^{L_{i}}\}$ (4)
As illustrated in Equation 5, we consider the hidden states $\mathbf{h}_{i}^{\{0\text{-}(L_{i}\text{-}1)\}}$ to derive the fused hidden states $\mathbf{h}_{k}^{\text{fused}}$. For the hidden states of each layer, a single scalar weight is assigned to control its importance. We utilize a set of scalar weight include $k$ elements $\{w_{i,j}^{1},\ldots,w_{i,j}^{k}\}$ for each hidden state $\mathbf{h}_{i,j}$ from all three encoders to generate three fused hidden states $\mathbf{h}_{\{1,\ldots,k\}}^{\text{fused}}$. Thus, the dimension of the whole matrices is $\mathbf{W}\in\mathbb{R}^{3\times(L_{1}+L_{2}+L_{3})}$.
$\mathbf{h}_{k}^{\text{fused}}=\sum_{i=1}^{3}\sum_{j=1}^{L_{i}}w_{i,j}^{k}\cdot\mathbf{h}_{i,j}$ (5)
Finally, we concatenate the last hidden states of the three encoders $\mathbf{h}_{\{1,2,3\}}^{L_{i}}$ with the three fused hidden states $\mathbf{h}_{\{1,\ldots,k\}}^{\text{fused}}$ along the feature dimension. Afterward, we apply a FFN to compress the feature dimension to match the dimension of the LLM embedding (Equation 2.2).
$\mathbf{h}^{\text{final}}$ $=\text{Concat}(\mathbf{h}_{\{1,2,3\}}^{L_{i}},\mathbf{h}_{\{1,\ldots,k\}}^{\text{fused}})$
@@ 2.3 Prompt Aware Routing
A prompt refers to a segment of text that provides context or objectives for the generation of the Speech LLM. Although the format of prompts can vary widely, they can typically be categorized into several distinct types according to the task the user wishes to perform. For instance, speech-related tasks, sound-related tasks and speech chat tasks Yang et al. (2024). In this paper, we investigate three tasks: ASR, speaker number verification, and AC. We utilize distinct experts for each task. For effective routing, the router must identify the task type based on the prompt. We employ a simple classification approach (Equation 7 and 8) wherein we use the last hidden states $\mathbf{H}_{\text{prompt}}$ of the prompt from the LLM, followed by a FFN and Softmax activation, to obtain the task posteriors $\text{P}(\text{Task}|\mathbf{H}_{\text{prompt}})$.
@@
To train the FFN, we create diverse prompts for each task using the LLM. Specifically, we manually write several prompts for each task and instruct ChatGPT to rewrite these prompts. Examples of these prompts are shown in Table 1. It is important to note that the audio features follow the prompt because we use the prompt to guide feature extraction and fusion. This approach differs from other works, where the audio features <|AUDIO|> can be positioned before the prompt.
ASR 1. Hear the audio clip and transform it into text format. <|AUDIO|>
2. Analyze the recording and summarize its contents. <|AUDIO|>
Table 1: Examples of prompts for different tasks.
@@ 2.4 Training Objective
The training loss function, as illustrated in Equation 2.4, is the sum of the cross-entropy loss $\mathcal{L}_{\text{G}}$ for prompt-aware gating and the cross-entropy loss $\mathcal{L}_{\text{llm}}$ between the LLM’s output $\mathbf{Y}$ and the ground truth $\hat{\mathbf{Y}}$.
@@ 3 Experimental Setups
In this section, we present a detailed description of our experimental setups, covering datasets, evaluation metrics, hyperparameters for the model architecture, and training. The links to the pretrained models and datasets used can be found in Appendix A.1, while the implementation details of baseline methods are available in Appendix A.2.
@@ 3.1 Datasets and Evaluation Metrics
We assess the efficacy of our method across three audio-to-text tasks: automatic speech recognition (ASR), speaker number verification (SNV), and audio captioning (AC). The training set includes the following: 200 hours of ASR data from LibriSpeech-100 Panayotov et al. (2015) and AMI Kraaij et al. (2005), 150 hours of SNV data synthesized from Common Voice V4 Ardila et al. (2019), and 100 hours of AC data from AudioCaps Kim et al. (2019). In total, the training data contains 450 hours of audio signals. For SNV, we randomly concatenate individual utterances to form new speech signals with the number of speakers ranging from one to four. The test dataset includes LibriSpeech-test-clean, LibriSpeech-test-other, AMI, the SNV test set from AIR-Bench Yang et al. (2024), and the test set of AudioCaps along with its corresponding QA version from AudioBench Wang et al. (2024), which contains diverse questions. ASR tasks focus on semantics. The LibriSpeech test set originates from audiobooks, demonstrating ASR performance in a clean scenario. AMI, a real meeting corpus containing spontaneous talk, reflects ASR performance in a more challenging, real-world scenario. SNV and AC test sets can indicate the Speech LLM’s ability to understand speaker and acoustic information. We evaluate the performance using word error rate (WER) for ASR tasks, accuracy for SNV, and METEOR Banerjee and Lavie (2005) for AC.
@@
   - WavLLM Hu et al. (2024) 4.95 (- 0.65) 09.19 (+0.27) 15.29 (+0.98) 39.20% (- 2.20) 34.93 (+1.97) 16.35 (+1.31) 2.67
   - SALMONN Tang et al. (2024) 5.04 (- 0.74) 09.70 (- 0.24) 19.04 (- 2.77) 49.40% (+8.00) 34.86 (+1.90) 15.97 (+0.93) 3.50
@@
PaM (Ours) 3.65 (+0.65) 07.07 (+2.39) 12.79 (+3.48) 42.80% (+1.40) 35.47 (+2.51) 15.70 (+0.66) 1.67
Table 2: Comparison of the proposed PaM method with single and multi-encoder baselines. For multi-encoder baselines: WavLLM and SALMONN use concatenation followed by a linear layer and Q-former, respectively, while ‘Average’ averages the three features. Values in the brackets indicate performance improvement (green) or degradation (red) compared to the best single encoder result. The avg rank column shows the average rank on each task. Smaller ranks indicate better performance.

### v2

@@ Abstract
Connecting audio encoders with large language models (LLMs) allows the LLM to perform various audio understanding tasks, such as automatic speech recognition (ASR) and audio captioning (AC). Most research focuses on training an adapter layer to generate a unified audio feature for the LLM. However, different tasks may require distinct features that emphasize either semantic or acoustic aspects, making task-specific audio features more desirable. In this paper, we propose Prompt-aware Mixture (PaM) to enhance the Speech LLM that uses multiple audio encoders. Our approach involves using different experts to extract different features based on the prompt that indicates different tasks. Experiments demonstrate that with PaM, only one Speech LLM surpasses the best performances achieved by all single-encoder Speech LLMs on ASR, Speaker Number Verification, and AC tasks. PaM also outperforms other feature fusion baselines, such as concatenation and averaging. Our code would be available at: https://github.com/shanweiqiao/PaM
@@
Weiqiao Shan1, Yuang Li2, Yuhao Zhang3, yingfeng luo1, Chen Xu4, Xiaofeng Zhao2, Long Meng1, Yunfei Lu2, Min Zhang2, Hao Yang2, Tong Xiao1,5thanks: Corresponding author, Jingbo Zhu1,5 1School of Computer Science and Engineering, Northeastern University, Shenyang, China 2Huawei Translation Services Center, Beijing, China 3The Chinese University of Hong Kong, Shenzhen, China 4College of Computer Science and Technology, Harbin Engineering University, Harbin, China 5NiuTrans Research, Shenyang, China
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance across various natural language processing tasks OpenAI (2023), paving the way for developing multimodal models Li et al. (2023); Xu et al. (2025); Wang et al. (2024b). In recent work, there has been a growing focus on merging speech encoders with LLMs, so that the LLM can understand the spoken content without the need for explicit transcription, promoting tasks such as direct speech translation Chen et al. (2024b) and named entity recognition from speech Li et al. (2024). Much of this work leverages adapter layers like attention layers Yu et al. (2024), adaptive CTC downsamplers Ling et al. (2023), and convolutional layers Fathullah et al. (2023) to downsample and map speech features into the LLM’s embedding space. Beyond semantic understanding tasks, Speech LLMs have been extended to encompass a broader range of applications, including audio event detection and audio captioning Chu et al. (2024).
Figure 1: ASR and Audio Caption tasks favor different encoders and layers of features. The x-axis corresponds to each layer of the encoder, while the bar chart illustrates the fine-grained layer importance, based on the normalized weight across layers from all encoders. The dotted lines indicate the average (AVG) importance of different encoders.
@@
Consequently, researchers have proposed using multiple encoders to extract more robust features. For instance, WavLLM Hu et al. (2024) employs both the WavLM Chen et al. (2022) and the Whisper Radford et al. (2022) encoder, while SALMONN Tang et al. (2024) integrates the Whisper encoder and the BEATs Chen et al. (2023). However, these approaches consider all encoders equally important and merge the features from different encoders based on a simple concatenation method across all tasks. As demonstrated in our experiments (Table 1), such conventional approaches can enhance performance in some tasks (e.g., audio captioning) but degrade others (e.g., ASR). Moreover, MoWE Zhang et al. (2024) employs a strong encoder and multiple weaker encoders via the Mixture of Experts (MoE) approach. However, MoWE only utilizes the input audio to control the routing mechanism, without incorporating task-specific information in prompts, which leads to suboptimal results (Table 11).
In this paper, we introduce Prompt-aware Mixture (PaM), a novel MoE method for merging multiple encoders to enhance Speech LLMs. Our approach integrates a prompt-aware routing mechanism, emphasizes feature fusion, and considers the relative importance of each encoder for different tasks, aiming to improve all downstream performance. PaM employs three distinct audio encoders: the Whisper encoder, WavLM, and Wav2Vec2 Baevski et al. (2020). We train a set of experts for prompt-aware feature fusion, comprising one shared expert and four task-specific experts. On each task, an expert learns the optimal weights for each encoder and its respective layers, and subsequently maps the resulting features to the embedding space of the Qwen2.5 model Team (2024). The embedding of the prompt is utilized to determine the appropriate routing. Notably, in PaM, the routing guides the selection of the fusion parameters rather than the choice of the encoder. Experiments are conducted across three tasks: ASR, speaker number verification(SNV), and AC. On all datasets, including LibriSpeech Panayotov et al. (2015), AMI Kraaij et al. (2005), AIR-Bench (SNV) Yang et al. (2024), and AudioCaps Kim et al. (2019), PaM achieves relative improvements of 15%, 25%, 3.4%, and 7.6%, respectively, in comparison to the best-performing single-encoder Speech LLM, and achieves a better average rank on all tasks compared with conventional approaches. Our contributions can be summarized as follows:
@@
  • We introduce PaM, a specialized MoE method that incorporates a prompt-aware routing mechanism to assign distinct weights to each encoder and its layers based on the task.
@@
Figure 2: The architecture of the proposed PaM method. The output feature $\mathbf{H}_{\text{prompt}}$ of the prompt $\mathbf{X}_{\text{prompt}}$ guides the routing of the MoE adapter, which incorporates a fixed shared expert (denoted as expert S) and a single routed expert (denoted as expert expert $i\in[1,N]$, where NN represents the total number of routed experts). For each expert, the last hidden states from all encoders $\mathbf{h}_{i}^{L_{i}-1}$ are concatenated with $K$ fused hidden states ($K$ being a hyperparameter, with the default value $K=3$) derived from a fusion weight matrix $\mathbf{W}$. Subsequently, a feedforward network (FFN) is applied to align with the dimensions of the LLM.
@@ 2.1 Overall Architecture
The architecture of the proposed PaM method is depicted in Figure 2 (left). As described in Equation 1, the LLM accepts the text prompt $\mathbf{X}_{\text{prompt}}$, which includes task-related information, along with the speech features $\mathbf{H}_{\text{audio}}$ as input, and subsequently generates the response $\mathbf{Y}$.
@@
To obtain $\mathbf{H}_{\text{audio}}$, we employ three encoders: the Whisper encoder, WavLM, and Wav2Vec2. For each encoder, the hidden states are initially processed by a feed-forward network (FFN) as described in Equation 2111The FFN module projects the hidden states from the encoder’s dimension to the LLM’s dimension, and maps the features from each encoder into a unified space that is shared across all encoders., resulting in the feature of each encoder, denoted as $\mathbf{H}_{i}$.
@@
$+\sum_{j=1}^{N}\text{G}_{j}(\mathbf{X}_{\text{prompt}})\times\text{Expert}_{j}(\mathbf{H}_{\{1,2,3\}})$ (3)
@@ 2.2 Multi-layer Fusion
We describe the multi-layer fusion process in Figure 2 (right). Different encoders exhibit distinct strengths. For example, WavLM is excellent at extracting speaker information Chen et al. (2022), while Wav2Vec2 excels in capturing semantic content Baevski et al. (2020). The Whisper encoder Radford et al. (2022), trained on a vast amount of data, provides superior features for AC and ASR in noisy environments [^2] These biases in the ability of different encoders on different tasks are also consistent with the results of our single-encoder baselines shown in Table 1. Additionally, features from different layers contain varying levels of information. Deeper layers hold high-level semantic information, whereas lower layers may contain fine-grained acoustic details. Thus, for feature fusion, we consider features from all layers of all three encoders. Specifically, for the feature $\mathbf{H}_{i}$ from a single encoder, it includes hidden states from all $L_{i}$ Transformer Vaswani et al. (2017) layers as well as $\mathbf{h}_{i}^{0}$, the hidden states following the convolutional layers (Equation 4).
$\mathbf{H}_{i}=\{\mathbf{h}_{i}^{0},\mathbf{h}_{i}^{1},...\mathbf{h}_{i}^{L_{i}-1}\}$ (4)
As illustrated in Equation 5, we consider the hidden states $\mathbf{h}_{i}^{\{0,1,\ldots,(L_{i}\text{-}2)\}}$ to derive the fused hidden states $\mathbf{h}_{k}^{\text{fused}}$. For each hidden state $\mathbf{h}^{l}_{i}$ of all three encoders, a set of scalar weights containing $\sum_{i}^{3}(L_{i}-1)$ elements is assigned to control its relative importance. We denote a set of scalar weights as $\{w_{i,k}^{l}|i\in[1,3],l\in[1,L_{i}-1]\}$ and then used it to generate fused hidden states $\mathbf{h}_{k}^{\text{fused}}$. To maintain the diversity of the fusion feature, we utilize $K$ sets of scalar weights. Thus, the dimension of the whole learnable matrices is $\mathbf{W}\in\mathbb{R}^{K\times(\sum_{i}^{3}(L_{i}-1))}$.
$\mathbf{h}_{k}^{\text{fused}}=\sum_{i=1}^{3}\sum_{l=1}^{L_{i}-1}w_{i,k}^{l}\cdot\mathbf{h}^{l}_{i}$ (5)
Finally, we concatenate the last hidden states of the three encoders $\mathcal{H}^{\text{last}}=\{\mathbf{h}_{i}^{L_{i}-1}|i\in[1,3]\}$ with the $K$ fused hidden states $\mathbf{h}_{\{1,\ldots,K\}}^{\text{fused}}$ along the feature dimension. Afterward, we apply an FFN to compress the feature dimension to match the dimension of the LLM embedding (Equation 2.2) [^3] We leverage both the final and fused hidden states, which are commonly used in semantic-related tasks (e.g., ASR) and acoustic-related tasks (e.g., audio captioning). The subsequent FFN module projects the concatenated features to the dimension of the LLM input embeddings, ensuring alignment between the feature space of speech features and text features..
$\mathbf{h}^{\text{final}}$ $=\text{Concat}(\mathcal{H}^{\text{last}},\mathbf{h}_{\{1,\ldots,K\}}^{\text{fused}})$
@@ 2.3 Prompt Aware Routing
A prompt refers to a text segment that provides context or objectives for generation, which can typically be categorized into several distinct types according to task. For instance, speech-related tasks, sound-related tasks, and speech chat tasks Yang et al. (2024). In this paper, we investigate three tasks: ASR, speaker number verification, and AC. We utilize distinct experts for each task. For effective routing, the router must identify the task type based on the prompt. We employ a simple classification approach (Equation 7 and 8) wherein we use the last hidden states $\mathbf{H}_{\text{prompt}}$ of the prompt from the LLM, followed by a FFN and Softmax activation, to obtain the task posteriors $\text{P}(\text{Task}|\mathbf{H}_{\text{prompt}})$.
@@
To train the FFN, we create diverse prompts for each task using the LLM. Specifically, we manually write several prompts for each task and instruct ChatGPT to rewrite these prompts. We list the examples of these prompts in Appendix A.1. It is important to note that the audio features follow the prompt because we use the prompt to guide feature extraction and fusion. This approach differs from other works, where the audio features <|AUDIO|> can be positioned before the prompt.
@@ 2.4 Training Objective
The training loss function, as illustrated in Equation 2.4, is the sum of the cross-entropy loss $\mathcal{L}_{\text{G}}$ for prompt-aware routing and the cross-entropy loss $\mathcal{L}_{\text{llm}}$ between the LLM’s output $\mathbf{Y}$ and the /ground truth $\hat{\mathbf{Y}}$. Loss function, as illustrated in Equation 2.4, is the sum of the cross-entropy loss $\mathcal{L}_{\text{G}}$ for prompt-aware routing and the cross-entropy loss $\mathcal{L}_{\text{llm}}$ between the LLM’s output $\mathbf{Y}$ and the ground truth $\hat{\mathbf{Y}}$.
@@ 3.1 Datasets and Evaluation Metrics
We assess the efficacy of our method across three audio-to-text tasks: automatic speech recognition (ASR), speaker number verification (SNV), and audio captioning (AC). We list the detailed information of training data in Appendix A.2. In total, the training data contains 450 hours of audio signals. The test dataset includes LibriSpeech-test-clean, LibriSpeech-test-other, AMI, the SNV test set from AIR-Bench, and the test set of AudioCaps along with its corresponding QA version from AudioBench Wang et al. (2024a), which contains diverse questions. ASR tasks focus on semantics. The LibriSpeech test set originates from audiobooks, demonstrating ASR performance in a clean scenario. AMI, a real meeting corpus containing spontaneous talk, reflects ASR performance in a more challenging, real-world scenario. SNV and AC test sets can indicate the Speech LLM’s ability to understand speaker and acoustic information. We evaluate the performance using word error rate (WER) for ASR tasks, accuracy for SNV, and METEOR Banerjee and Lavie (2005) for AC. Additionally, we list the results on AC and AC QA tasks with more metrics in Appendix A.4.
@@
   - WavLLM Hu et al. (2024) 4.95 (- 0.65) 9.19 (+0.27) 15.29 (+0.98) 39.20% (- 2.20) 34.93 (+1.97) 16.35 (+1.31) 2.67
   - SALMONN Tang et al. (2024) 5.04 (- 0.74) 9.70 (- 0.24) 19.04 (- 2.77) 49.40% (+8.00) 34.86 (+1.90) 15.97 (+0.93) 3.50
@@
PaM (Ours) 3.65 (+0.65) 7.07 (+2.39) 12.79 (+3.48) 42.80% (+1.40) 35.47 (+2.51) 15.70 (+0.66) 1.67
Table 1: Comparison of the proposed PaM method with single and multi-encoder baselines. Values in the brackets indicate performance improvement (green) or degradation (red) compared to the best single encoder result. The AVG rank column shows the average rank on each task. Smaller ranks indicate better performance.